# Analytically Continuing the Randomized Measurement Toolbox

### v2

@@ Abstract
Introduction.– Recent advances in quantum simulator platforms enable the study of complex quantum many-body phenomena, such as thermalization, scrambling, and topological order, in well-controlled laboratory settings [Preskill2018NISQ, Altman2021PRXQReview]. Understanding these phenomena requires access to the entanglement structure of quantum states. Quantum information theory provides a quantitative framework for characterizing entanglement through the von Neumann entanglement entropy of a density matrix $\rho$,
@@
which serves as a fundamental measure of bipartite entanglement [horodecki2009quantum]. It plays a central role in diagnosing topological order [TEE_KitaevPreskill, TEE_LevinWen], probing thermalization and information scrambling [Fast_Scramblers, ETH_EE, Deutsch_2018, Kaufman_2016, EntanglementGrowthRUCs], and studying quantum criticality [EE_Criticality, EE_PhaseTransition] and the emergence of bulk gravity in AdS/CFT [Ryu_2006, Ryu_2007]. However, directly measuring the von Neumann entropy for large quantum systems in quantum simulators remains experimentally challenging, motivating the search for scalable measurement approaches.
@@
While full quantum state tomography allows complete reconstruction of a quantum state and thereby access to its von Neumann entropy, its sample complexity scales sharply exponentially with system size, rendering it impractical beyond few-qubit systems [haah2017sample, ODonnell2016Jun]. Recent advances in randomized measurement (RM) protocols [elben2019statistical, huang2020predicting, elben2023randomized] have pushed the boundary to substantially larger system sizes for a broad class of polynomial functions of the density matrix [vitale2024robust]. Such functions can also be estimated utilizing many-body interference between multiple physical copies of quantum state prepared simultaneously in a quantum experiment [islam2015measuring, kaufman2016]. These protocols enable measurement of scalable quantities such as Rényi entropies $S_{k}(\rho)=\frac{1}{1-k}\log_{2}\mathrm{Tr}\rho^{k}$, $k\geq 2$ integer [elben2018renyi, elben2019statistical], Hilbert–Schmidt fidelities, and mixed-state entanglement witnesses [elben2020mixed]. Such quantities serve as powerful diagnostics of correlations and entanglement [brydges2019probing, elben2020mixed], but they are not entanglement measures [horodecki2009quantum]. Moreover, Rényi entropies can display qualitatively different behavior from the von Neumann entropy [tibor2019diffusion, wang2025singularity, bertini2022growth, agon2023renyi]. Therefore, reliable estimates of the von Neumann entropy are essential for a quantitative characterization of entanglement in many-body quantum systems.
@@
Figure 1: SAC and traditional fitting methods [vermersch2024enhanced, ChebyshevBook] estimating the half-chain von Neumann entropy in the ground state of the transverse-field Ising model $H=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-h\sum_{i}\sigma_{i}^{x}$ away from criticality ($J=1$, $h=0.5$, 15 sites). Error percentage of different methods are shown as a function of the largest Rényi index in the dataset. Left: for noiseless Rényi inputs. Right: for inputs with 10% independent Gaussian noise, averaged over 200 realizations.
@@
In this Letter, we develop a framework to reliably estimate the von Neumann entropy in quantum simulation experiments using a finite number of measured Rényi entropies. While the von Neumann entropy can, in principle, be obtained from knowledge of all integer Rényi entropies via analytic continuation, $S_{\mathrm{vN}}=\lim_{k\to 1^{+}}S_{k}$, a consequence of Carlson’s theorem [hardy1920two], in practice only a limited number of Rényi orders are accessible, each with finite statistical uncertainty. This renders the continuation problem ill-posed and highly sensitive to statistical noise—much like the analytic continuation of Green’s functions away from the Matsubara domain or extracting physical scattering amplitudes in quantum chromodynamics [Maximum_Entropy_Method, Bayesian_Inference, SVZ_SumRules, Sum_Rules_Review, Smearing_Method]. We employ the stabilized analytic continuation (SAC) approach of Ciulli and Spearman [CS1, CS2, CS3, CS4, CS5], which reformulates analytic continuation as a constrained convex optimization problem: among all analytic functions compatible with the data, SAC selects the one minimizing a chosen $L^{2}$-norm. Our novel adaption of SAC yields stable and noise-resilient estimates of the von Neumann entropy even when only a few Rényi entropies are available. Fig. 1 shows the performance of SAC in an example of estimating the half-chain von Neumann entropy of a spin chain model, and compares it against more traditional extrapolation schemes. For benchmarking, we synthetically add independent Gaussian noise to each input Rényi entropies. Beyond entanglement entropy, our framework applies broadly to estimating general non-polynomial spectral functions of quantum states accessible through analytic continuation.
@@
In the following, we present our adaptation of the SAC method, including deriving an explicit analytic expression of the von Neumann entropy estimate from a few Rényi entropies, and generalizing to the case with (correlated) statistical noise. We benchmark our framework using numerically simulated Rényi data for a quenched Néel state and demonstrate its performance on experimental data from a trapped-ion quantum simulator [brydges2019probing], where it enables extraction of subsystem von Neumann entropies across time.
@@
Stabilized analytic continuation.– Let $S_{z}(\rho)=\frac{1}{1-z}\log_{2}\mathrm{Tr}\rho^{z}$, $z\in\mathbb{C}$, denote the Rényi function on the complex plane. Suppose we are provided noisy estimates of this function at $N$ integer points $z=2,3,.\cdots,k_{\text{max}}$. Our task is to analytically continue this dataset to the von Neumann point, $z=1$. Assuming the true underlying function is smooth, traditional approaches typically involve fitting with lower-degree polynomials to mitigate overfitting [StatLearningBook, mohammadipour2025ZNE]. Nevertheless, choosing a low-degree polynomial still requires model assumptions and entails an arbitrary selection of polynomial basis.
@@
This leads us to the method of stabilized analytic continuation (SAC), which bypasses these challenges by exploiting the fundamental analytic properties of the underlying function [CS1, CS2, CS3, CS4, CS5]. To adapt it to the von Neumann entropy problem, we construct a variational function whose minimum $L^{2}$-norm determines the optimal estimate of the von Neumann entropy. We explicitly derive the analytical expression for the von Neumann estimate resulting from this variational approach in the noiseless case. We also extend the SAC framework, originally formulated for uncorrelated noise, to handle correlated noise. Our adaptation is found to be more robust to realistic noise than conventional extrapolation techniques, as confirmed using data with correlated noise below.
@@
    For fixed $\alpha$, we search for the analytic function $Y_{\alpha}(w)$ which takes the desired values $\{Y_{\alpha}(w_{i})=D^{\prime}_{\alpha}(w_{i})=d_{i}(\alpha)-d_{2}(\alpha)\}$ at the points $\{w_{i}\}$ and minimizes the norm $Y_{\alpha}||$. This minimization can be recast into a linear matrix optimization problem by leveraging the analyticity of $Y_{\alpha}(w)$ [CS1]. The minimal norm
@@
In the presence of noise, Step 1 above is modified to include sampling of all data points that lie within a suitable neighbourhood of the mean values $D^{\prime}_{\alpha}(w_{j})$. If $C$ denotes the covariance matrix of the data points, then the goodness of fit of an arbitrary point $\{y_{j}\}$ in data space can be measured using a $\chi^{2}$ statistic: $\chi^{2}(\mathbf{y};\alpha)=\sum_{i,j=2}^{k_{\text{max}}}(y_{i}-D^{\prime}_{\alpha}(w_{i}))(C^{\prime-1})_{ij}(y_{j}-D^{\prime}_{\alpha}(w_{j}))$, where $\mathbf{y}=\{y_{2},...,y_{k_{\text{max}}}\}$, and $C^{\prime}_{ij}=\frac{C_{ij}}{(z_{i}-1)(z_{j}-1)}$ is the covariance matrix of the corresponding discrepancy function. Note that for the noisy data case, we treat the subtraction point as a variational parameter as well [ciulli1982analytic, ciulli1983search], and hence the sum runs from $i,j=2$. The minimal norm $\delta(\alpha)$ is now computed by minimizing over all possible points y subject to the constraint $\chi^{2}(\mathbf{y};\alpha)\leq\chi^{2}_{0}$ for some constant $\chi^{2}_{0}$ typically chosen to be $\mathcal{O}(k_{\mbox{\tiny max}})$. Namely,
@@
Randomized measurements and classical shadow tomography.— We briefly summarize the randomized measurement (RM) protocol [elben2019statistical, huang2020predicting, elben2023randomized]. For a system of $L$ qubits in an unknown state $\rho$, local random unitaries $U=U_{1}\otimes\cdots\otimes U_{L}$ are drawn independently from a unitary 3-design [elben2023randomized] and followed by projective measurements in the computational basis. From the resulting bit strings, collected over $N_{u}$ unitaries and $N_{m}$ repetitions, one constructs classical shadows $\rho^{(m)}$}, which are unbiased estimators of the density matrix, $\mathbb{E}[\rho^{(m)}]=\rho$. Expectation values of observables and polynomial functions of $\rho$, such as trace moments $p_{k}=\mathrm{Tr}(\rho^{k})$, can then be estimated via U-statistics [elben2020mixed]:
@@
where $I_{k}$ denotes distinct index sets of size $k$. Because evaluating $\hat{p}_{k}$ scales as $\mathcal{O}(N_{u}^{k})$, implying significant post-processing time, the batch-shadow variant [rathentanglement2023] groups $N_{u}$ shadows into $N_{B}\ll N_{u}$ batches, averaging within each batch before computing the U-statistic. This strongly reduces the computational cost while retaining near-optimal statistical accuracy [rathentanglement2023]; we employ this variant throughout, typically with $N_{B}=10\sim 20$.
While in principle Rényi entropies up to order $N_{u}$ ($N_{B}$) could be estimated from the data, higher-order Rényi entropies suffer from rapidly increasing statistical uncertainty [elben2020mixed], yielding in practice a finite set of accessible Rényi orders. In addition, shadow tomography cannot directly estimate non-polynomial functionals such as the von Neumann entropy. In the following, we show that SAC allows accurate inference from finite-order Rényi entropies (up to $k_{\text{max}}=6$) obtained via shadow tomography, yielding robust von Neumann entropy estimates for both simulated and experimental data.
Benchmarking SAC with numerical simulations.— We demonstrate the performance of the SAC method on synthetic data generated from numerical simulations of quench dynamics in a 10-qubit system. The system is initialized in a low-entropy Néel state and subsequently evolved under a spin-$1/2$ Hamiltonian [brydges2019probing, Porras2004PRL],
@@
where $\sigma_{i}^{\pm}$ are spin ladder operators and the couplings follow an approximate power-law decay $J_{ij}\approx J/|i-j|^{1.2}$ with $J=420\mathrm{s}^{-1}\ll B$. The simulations incorporate realistic decoherence sources, including imperfect Néel-state preparation, spontaneous emission, spin flips, and depolarization from local unitaries [brydges2019probing].
RM are simulated on a classical computer to construct classical (batch) shadows, which are then used to estimate Rényi entropies [elben2020mixed]. We perform 1000 independent shadow experiments, each consisting of $N_{u}=500$ random unitaries and $N_{m}=150$ measurements per unitary. Using U-statistics, Rényi entropies of orders $k=2,\cdots,6$ are computed for each experiment. The von Neumann entropy is extracted with both SAC and traditional fitting schemes, including Chebyshev polynomial interpolation [ChebyshevBook] and least-squares fitting [vermersch2024enhanced]. These estimates are compared with the exact von Neumann entropy obtained directly from the simulation (Fig. 2). For SAC, the 1000 Rényi estimates are grouped into 50 groups of 20 Rényi sets each. For each group, we calculate the mean and covariance matrix of the 20 Rényi sets and feed these to the noisy SAC protocol. The resulting 50 von Neumann entropy estimates are then averaged to obtain the final mean and statistical uncertainty. Results across different subsystem sizes and timesteps are summarized in Fig. 2.
@@
Discussion.– We introduced a robust framework for extracting the von Neumann entropy from finite sets of Rényi entropies subject to statistical noise obtained via randomized measurements. When tested on extensive simulation data, our method outperforms conventional fitting approaches. We applied our method to experimental data from a trapped-ion quantum simulator, obtaining direct estimates of subsystem von Neumann entropy growth following a quantum quench. Since randomized measurements are now routinely used to measure entropies in many platforms, our noise-resilient approach provides a practical post-processing tool for analyzing data affected by statistical uncertainty and may substantially reduce measurement overheads. It also extends to situations where Renyi entropies are measured through many-body interference experiments [islam2015measuring, kaufman2016].
There are many potential applications of SAC beyond extracting von Neumann entropies. The method naturally extends to any information-theoretic quantity accessible through a replica trick—such as logarithmic negativity, Rényi relative entropies, Petz–Rényi divergences, or Uhlmann fidelity [elben2020mixed, lieb1973convex, uhlmann1977relative, petz1986quasi, Lennert2013RenyiEntropies, Wilde2014sandwiched, Datta2015alphaz, KudlerFlam2024RMI]—which we leave for future exploration. More broadly, SAC offers a promising strategy for addressing long-standing analytic-continuation challenges in many other areas of physics, from quantum Monte Carlo studies to the reconstruction of scattering amplitudes.
@@
Note: The open-source code to extract Rényi entropies from batch shadows, applied to the experimental data of Ref. [brydges2019probing], is available on Github, and presented in details in Ref. [elben2025randommeasjl].

### v3

@@ Abstract
Introduction.– Recent advances in quantum simulator platforms enable the study of complex quantum many-body phenomena, such as thermalization, scrambling, and topological order, in well-controlled laboratory settings [44, 2]. Understanding these phenomena requires access to the entanglement structure of quantum states. Quantum information theory provides a quantitative framework for characterizing entanglement through the von Neumann entanglement entropy of a density matrix $\rho$,
@@
which serves as a fundamental measure of bipartite entanglement [25]. It plays a central role in diagnosing topological order [31, 33], probing thermalization and information scrambling [49, 19, 13, 30, 38], and studying quantum criticality [53, 40] and the emergence of bulk gravity in AdS/CFT [48, 47]. However, directly measuring the von Neumann entropy for large quantum systems in quantum simulators remains experimentally challenging, motivating the search for scalable measurement approaches.
@@
While full quantum state tomography allows complete reconstruction of a quantum state and thereby access to its von Neumann entropy, its sample complexity scales sharply exponentially with system size, rendering it impractical beyond few-qubit systems [22, 39]. Recent advances in randomized measurement (RM) protocols [17, 26, 14] have pushed the boundary to substantially larger system sizes for a broad class of polynomial functions of the density matrix [54]. Such functions can also be estimated utilizing many-body interference between multiple physical copies of quantum state prepared simultaneously in a quantum experiment [27, 29]. These protocols enable measurement of scalable quantities such as Rényi entropies $S_{k}(\rho)=\frac{1}{1-k}\log_{2}\mathrm{Tr}\rho^{k}$, $k\geq 2$ integer [16, 17], Hilbert–Schmidt fidelities, and mixed-state entanglement witnesses [15]. Such quantities serve as powerful diagnostics of correlations and entanglement [5, 15], but they are not entanglement measures [25]. Moreover, Rényi entropies can display qualitatively different behavior from the von Neumann entropy [45, 55, 4, 1]. Therefore, reliable estimates of the von Neumann entropy are essential for a quantitative characterization of entanglement in many-body quantum systems.
@@
Figure 1: SAC and traditional fitting methods [52, 35] estimating the half-chain von Neumann entropy in the ground state of the transverse-field Ising model $H=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-h\sum_{i}\sigma_{i}^{x}$ away from criticality ($J=1$, $h=0.5$, 15 sites). Error percentage of different methods are shown as a function of the largest Rényi index in the dataset. Left: for noiseless Rényi inputs. Right: for inputs with 10% independent Gaussian noise, averaged over 200 realizations.
@@
In this Letter, we develop a framework to reliably estimate the von Neumann entropy in quantum simulation experiments using a finite number of measured Rényi entropies. While the von Neumann entropy can, in principle, be obtained from knowledge of all integer Rényi entropies via analytic continuation, $S_{\mathrm{vN}}=\lim_{k\to 1^{+}}S_{k}$, a consequence of Carlson’s theorem [23], in practice only a limited number of Rényi orders are accessible, each with finite statistical uncertainty. This renders the continuation problem ill-posed and highly sensitive to statistical noise—much like the analytic continuation of Green’s functions away from the Matsubara domain or extracting physical scattering amplitudes in quantum chromodynamics [21, 28, 50, 12, 42]. We employ the stabilized analytic continuation (SAC) approach of Ciulli and Spearman [8, 9, 6, 7, 20], which reformulates analytic continuation as a constrained convex optimization problem: among all analytic functions compatible with the data, SAC selects the one minimizing a chosen $L^{2}$-norm. Our novel adaption of SAC yields stable and noise-resilient estimates of the von Neumann entropy even when only a few Rényi entropies are available. Fig. 1 shows the performance of SAC in an example of estimating the half-chain von Neumann entropy of a spin chain model, and compares it against more traditional extrapolation schemes. For benchmarking, we synthetically add independent Gaussian noise to each input Rényi entropies. Beyond entanglement entropy, our framework applies broadly to estimating general non-polynomial spectral functions of quantum states accessible through analytic continuation.
@@
In the following, we present our adaptation of the SAC method, including deriving an explicit analytic expression of the von Neumann entropy estimate from a few Rényi entropies, and generalizing to the case with (correlated) statistical noise. We benchmark our framework using numerically simulated Rényi data for a quenched Néel state and demonstrate its performance on experimental data from a trapped-ion quantum simulator [5], where it enables extraction of subsystem von Neumann entropies across time.
@@
Stabilized analytic continuation.– Let $S_{z}(\rho)=\frac{1}{1-z}\log_{2}\mathrm{Tr}\rho^{z}$, $z\in\mathbb{C}$, denote the Rényi function on the complex plane. Suppose we are provided noisy estimates of this function at $N$ integer points $z=2,3,.\cdots,k_{\text{max}}$. Our task is to analytically continue this dataset to the von Neumann point, $z=1$. Assuming the true underlying function is smooth, traditional approaches typically involve fitting with lower-degree polynomials to mitigate overfitting [24, 36]. Nevertheless, choosing a low-degree polynomial still requires model assumptions and entails an arbitrary selection of polynomial basis.
@@
This leads us to the method of stabilized analytic continuation (SAC), which bypasses these challenges by exploiting the fundamental analytic properties of the underlying function [8, 9, 6, 7, 20]. To adapt it to the von Neumann entropy problem, we construct a variational function whose minimum $L^{2}$-norm determines the optimal estimate of the von Neumann entropy. We explicitly derive the analytical expression for the von Neumann estimate resulting from this variational approach in the noiseless case. We also extend the SAC framework, originally formulated for uncorrelated noise, to handle correlated noise. Our adaptation is found to be more robust to realistic noise than conventional extrapolation techniques, as confirmed using data with correlated noise below.
@@
    For fixed $\alpha$, we search for the analytic function $Y_{\alpha}(w)$ which takes the desired values $\{Y_{\alpha}(w_{i})=D^{\prime}_{\alpha}(w_{i})=d_{i}(\alpha)-d_{2}(\alpha)\}$ at the points $\{w_{i}\}$ and minimizes the norm $Y_{\alpha}||$. This minimization can be recast into a linear matrix optimization problem by leveraging the analyticity of $Y_{\alpha}(w)$ [8]. The minimal norm
@@
In the presence of noise, Step 1 above is modified to include sampling of all data points that lie within a suitable neighbourhood of the mean values $D^{\prime}_{\alpha}(w_{j})$. If $C$ denotes the covariance matrix of the data points, then the goodness of fit of an arbitrary point $\{y_{j}\}$ in data space can be measured using a $\chi^{2}$ statistic: $\chi^{2}(\mathbf{y};\alpha)=\sum_{i,j=2}^{k_{\text{max}}}(y_{i}-D^{\prime}_{\alpha}(w_{i}))(C^{\prime-1})_{ij}(y_{j}-D^{\prime}_{\alpha}(w_{j}))$, where $\mathbf{y}=\{y_{2},...,y_{k_{\text{max}}}\}$, and $C^{\prime}_{ij}=\frac{C_{ij}}{(z_{i}-1)(z_{j}-1)}$ is the covariance matrix of the corresponding discrepancy function. Note that for the noisy data case, we treat the subtraction point as a variational parameter as well [10, 11], and hence the sum runs from $i,j=2$. The minimal norm $\delta(\alpha)$ is now computed by minimizing over all possible points y subject to the constraint $\chi^{2}(\mathbf{y};\alpha)\leq\chi^{2}_{0}$ for some constant $\chi^{2}_{0}$ typically chosen to be $\mathcal{O}(k_{\mbox{\tiny max}})$. Namely,
@@
Randomized measurements and classical shadow tomography.— We briefly summarize the randomized measurement (RM) protocol [17, 26, 14]. For a system of $L$ qubits in an unknown state $\rho$, local random unitaries $U=U_{1}\otimes\cdots\otimes U_{L}$ are drawn independently from a unitary 3-design [14] and followed by projective measurements in the computational basis. From the resulting bit strings, collected over $N_{u}$ unitaries and $N_{m}$ repetitions, one constructs classical shadows $\rho^{(m)}$}, which are unbiased estimators of the density matrix, $\mathbb{E}[\rho^{(m)}]=\rho$. Expectation values of observables and polynomial functions of $\rho$, such as trace moments $p_{k}=\mathrm{Tr}(\rho^{k})$, can then be estimated via U-statistics [15]:
@@
where $I_{k}$ denotes distinct index sets of size $k$. Because evaluating $\hat{p}_{k}$ scales as $\mathcal{O}(N_{u}^{k})$, implying significant post-processing time, the batch-shadow variant [46] groups $N_{u}$ shadows into $N_{B}\ll N_{u}$ batches, averaging within each batch before computing the U-statistic. This strongly reduces the computational cost while retaining near-optimal statistical accuracy [46]; we employ this variant throughout, typically with $N_{B}=10\sim 20$.
While in principle Rényi entropies up to order $N_{u}$ ($N_{B}$) could be estimated from the data, higher-order Rényi entropies suffer from rapidly increasing statistical uncertainty [15], yielding in practice a finite set of accessible Rényi orders. In addition, shadow tomography cannot directly estimate non-polynomial functionals such as the von Neumann entropy. In the following, we show that SAC allows accurate inference from finite-order Rényi entropies (up to $k_{\text{max}}=6$) obtained via shadow tomography, yielding robust von Neumann entropy estimates for both simulated and experimental data.
Benchmarking SAC with numerical simulations.— We demonstrate the performance of the SAC method on synthetic data generated from numerical simulations of quench dynamics in a 10-qubit system. The system is initialized in a low-entropy Néel state and subsequently evolved under a spin-$1/2$ Hamiltonian [5, 43],
@@
where $\sigma_{i}^{\pm}$ are spin ladder operators and the couplings follow an approximate power-law decay $J_{ij}\approx J/|i-j|^{1.2}$ with $J=420\mathrm{s}^{-1}\ll B$. The simulations incorporate realistic decoherence sources, including imperfect Néel-state preparation, spontaneous emission, spin flips, and depolarization from local unitaries [5].
RM are simulated on a classical computer to construct classical (batch) shadows, which are then used to estimate Rényi entropies [15]. We perform 1000 independent shadow experiments, each consisting of $N_{u}=500$ random unitaries and $N_{m}=150$ measurements per unitary. Using U-statistics, Rényi entropies of orders $k=2,\cdots,6$ are computed for each experiment. The von Neumann entropy is extracted with both SAC and traditional fitting schemes, including Chebyshev polynomial interpolation [35] and least-squares fitting [52]. These estimates are compared with the exact von Neumann entropy obtained directly from the simulation (Fig. 2). For SAC, the 1000 Rényi estimates are grouped into 50 groups of 20 Rényi sets each. For each group, we calculate the mean and covariance matrix of the 20 Rényi sets and feed these to the noisy SAC protocol. The resulting 50 von Neumann entropy estimates are then averaged to obtain the final mean and statistical uncertainty. Results across different subsystem sizes and timesteps are summarized in Fig. 2.
@@
Discussion.– We introduced a robust framework for extracting the von Neumann entropy from finite sets of Rényi entropies subject to statistical noise obtained via randomized measurements. When tested on extensive simulation data, our method outperforms conventional fitting approaches. We applied our method to experimental data from a trapped-ion quantum simulator, obtaining direct estimates of subsystem von Neumann entropy growth following a quantum quench. Since randomized measurements are now routinely used to measure entropies in many platforms, our noise-resilient approach provides a practical post-processing tool for analyzing data affected by statistical uncertainty and may substantially reduce measurement overheads. It also extends to situations where Renyi entropies are measured through many-body interference experiments [27, 29].
There are many potential applications of SAC beyond extracting von Neumann entropies. The method naturally extends to any information-theoretic quantity accessible through a replica trick—such as logarithmic negativity, Rényi relative entropies, Petz–Rényi divergences, or Uhlmann fidelity [15, 34, 51, 41, 37, 56, 3, 32]—which we leave for future exploration. More broadly, SAC offers a promising strategy for addressing long-standing analytic-continuation challenges in many other areas of physics, from quantum Monte Carlo studies to the reconstruction of scattering amplitudes.
@@
Note: The open-source code to extract Rényi entropies from batch shadows, applied to the experimental data of Ref. [5], is available on Github, and presented in details in Ref. [18].